### v2

# MyFramelinecolor=black, outerlinewidth=2pt, roundcorner=20pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=codeBackground

Abstract
We revisit the longstanding open problem of implementing Nakamoto’s proof-of-work (PoW) consensus based on a real-world computational task $T(x)$ (as opposed to artificial random hashing), in a truly permissionless setting where the miner itself chooses the input $x$. The challenge in designing such a Proof-of-Useful-Work (PoUW) protocol, is using the native computation of $T(x)$ to produce a PoW certificate with prescribed hardness and with negligible computational overhead over the worst-case complexity of $T(\cdot)$ – This ensures malicious miners cannot “game the system" by fooling the verifier to accept with higher probability compared to honest miners (while using similar computational resources). Indeed, obtaining a PoUW with $O(1)$-factor overhead is trivial for any task $T$, but also useless.
Our main result is a PoUW for the task of Matrix Multiplication $\mathsf{MatMul}(A,B)$ of arbitrary matrices with $1+o(1)$ multiplicative overhead compared to naïve $\mathsf{MatMul}$ (even in the presence of Fast Matrix Multiplication-style algorithms, which are currently impractical). We conjecture that our protocol has optimal security in the sense that a malicious prover cannot obtain any significant advantage over an honest prover. This conjecture is based on reducing hardness of our protocol to the task of solving a batch of low-rank random linear equations which is of independent interest.
Since $\mathsf{MatMul}$s are the bottleneck of AI compute as well as countless industry-scale applications, this primitive suggests a concrete design of a new L1 base-layer protocol, which nearly eliminates the energy-waste of Bitcoin mining – allowing GPU consumers to reduce their AI training and inference costs by “re-using" it for blockchain consensus, in exchange for block rewards (2-for-1). This blockchain is currently under construction.
1 Introduction
The concept of proofs of work (PoW), i.e., easy-to-check proofs of computational effort, was proposed in 1992 in a seminal work of Dwork and Naor [DN92]. Their original motivation was to discourage spam email messages by requiring the sender to compute a moderately hard, but not intractable, function in order for the email to go through. This idea is far more general than combating spam emails; indeed, it can be used as a method for limiting access to any resource.
[…]
Perhaps the most well-known application of Back’s PoW idea is in the Bitcoin cryptocurrency network [Nak08, GKL24]. Instead of using PoW to deter malicious email senders, Bitcoin’s PoW is used to enable competitive mining and thereby to secure the network. Specifically, a Bitcoin miner runs a program that collects unconfirmed transactions from the network to form a block. A block is only accepted by the network if its hash meets a certain difficulty target, (essentially) described by the number of trailing 0s needed in the hash output.
PoW Criticism.
Undoubtedly, the realization of the Bitcoin network and its underlying technology is a revolutionary achievement in distributed computing, with far-reaching applications in both theory and practice. Notwithstanding, a significant setback and common point of criticism on the technology is its energy-intensive nature and exorbitant environmental costs (consuming over 2% annual electricity in the US alone). [^2] https://ccaf.io/cbnsi/cbeci/ghg/comparisons (accessed 02.04.2025). To make matters worse, the need for continual acquisition of specialized hardware (ASICs) to retain competitive mining, proliferates electronic waste generation, as these devices quickly become obsolete.
Thus, an outstanding question in this context is whether it is necessary for the mining process to be so wasteful and have no other use beyond securing and driving the blockchain. A similar question was raised already in the 1992 work of Dwork and Naor, in the context of using PoW as a spam filter [DN92]:
“Finally, the evaluation of the pricing function serves no useful purpose, except serving as a deterrent. It would be exciting to come up with a scheme in which evaluating the pricing function serves some additional purpose.”
Proofs of Useful Work.
The above challenge is nowadays commonly referred to as a proof of useful work (PoUW). Intuitively, a proof of work is useful if it simultaneously satisfies the following properties:
  • Economic value: The results of computations performed by miners should be of interest, and moreover, someone must be willing to pay for the result of the computation (formally, a task is $\kappa$-useful, if someone is willing to pay $\kappa\$$ for its completion).
  • Efficiency: The algorithm used to solve the problems is essentially optimal. This property is necessary to make sure that there is no “waste” in the mining process as compared to an external setting.
  • Security: It should be infeasible to fool the verifier into believing a problem was solved using bounded resources while it was not.
The first formal treatment of PoUW we are aware of is due to Ball et al. [BRSV17] wherein the authors suggest a proof of work mechanism (in the plain model) based on the conjectured hardness of various fine-grained algorithmic problems. They termed the proof of work useful because the miner could choose the instance of the task it wishes to solve by itself. Unfortunately, the overhead of the honest prover was significant, requiring the prover to solve poly-log many different instances of the task just to convince the verifier that a single instance was solved. In other words, their scheme did not satisfy the efficiency property from above and in a follow-up version of their manuscript [BRSV18] they retracted all claims about “usefulness.”
The absence of a PoUW (prior to this work) is not due to lack of trying. Not only was this problem mentioned by Dwork and Naor [DN92], PoUW has been the holy-grail of distributed consensus since the early days of Bitcoin, because it avoids the tradeoff between resource-efficiency and security, present in proof-of-stake (PoS) or other energy-efficient alternatives. As Vitalik Buterin, co-founder of Ethereum, articulates in his 2019 blog post: [^3] https://vitalik.eth.limo/general/2019/11/22/progress.html, Bullet 7.
“If either an efficiently verifiable proof-of-computation for Folding@home can be produced, or if we can find some other useful computation which is easy to verify, then cryptocurrency mining could actually become a huge boon to society, not only removing the objection that Bitcoin wastes "energy", but even being socially beneficial by providing a public good."
In the same post, Buterin adds that the challenge of obtaining a proof of useful work is “probably not feasible”. The transition of Etherium from PoW to PoS is, to a large extent, attributed to this impossibility conjecture.
Even a theoretical construction of PoUW was elusive. Specifically, we are not aware of any suggestion for a truly useful proof of work mechanism where the miner (and not the network) gets to choose their own instance. We refer to a manuscript of Dotan and Tochner [DT20] surveying prior approaches, identifying for each one which properties of a “non-wasteful” and “truly useful” proof of work they do not satisfy.
1.1 Our Contributions
Our main contribution is the first construction of a truly useful PoW protocol, for the omnipresent task of matrix multiplication (MatMul). At a high level, our (GPU-compatible) protocol can re-use native $\textsc{MatMul}(A,B)$ computations of arbitrary matrices (submitted by the miner at her own discretion/application), to produce a verifiable PoW certificate obeying Nakamoto’s Poisson distribution, with $1+o(1)$ multiplicative overhead over the worst-case time for computing $A\cdot B$ (e.g., $(1+o(1))n^{3}$ in the case of square-matrices, assuming naïve MatMul is the baseline).
Before we describe the protocol (called cuPOW; see Section 2), let us first review the axioms of PoUW, and argue why the MatMul primitive satisfies them.
  • Economic value: Matrix multiplication is the backbone of many industry-scale applications. In particular, they are (by far) the bottleneck of AI training and inference, where giant MatMuls (as large as 40K x 40K in LLMs) are required for both forward and backward propagation. In short, MatMuls are the primary reason for the exorbitant electricity costs of AI and for its adoption barrier by millions of businesses (c.f. GPT-4 training estimated at \$300 million and the next model of OpenAI is projected to cost 100x more). See Section 7 for an overview. Beyond AI, matrix multiplication is a central operation in many other applications, including vector databases, quantum simulations, graphics rendering and statistical physics to mention a few.
  • Efficiency: Our method for matrix multiply has minimal overhead over the direct matrix multiplication algorithm. Specifically, the running time of our algorithm is $N\cdot(1+o(1))$ where $N$ is a well-known baseline implementation of matrix multiplication (in practice, $N=O(n^{3})$). We do not introduce any hidden large constant or significant terms. Thus, our protocol is not only theoretically interesting, but is also very promising from a practical perspective.
  • Security: Our protocol’s soundness relies on a hardness assumption of a certain direct-product problem of correlated instances (essentially random low-rank linear equations). Under this direct product conjecture, no adversary with comparable runtime can obtain non-trivial advantage which causes the verifier to accept with higher probability. In particular, multiplying the all-zeroes matrices $A=B={\bf 0\rm}^{n\times n}$ using our protocol is as hard as any other product!
MatMul algorithms.
The discovery of Fast Matrix Multiplication (FMM) algorithms by Strassen [Str69] and subsequent improvements (starting with [CW87] and most recently [VXXZ24]), allowing to multiply $n\times n$ matrices in sub-cubic time $n^{\omega}\ll n^{3}$, had a profound theoretical impact in computer science and algorithm design. Our PoW protocol can be instantiated and applied with any of the above FMMs as the baseline algorithm. Unfortunately, FMM algorithms are unlikely to be practical on any imaginable hardware, as they are are highly asymptotic (earning them the infamous name “galactic algorithms"), or require significant memory.
Our protocol’s description and security analysis hold even in the presence of FMM algorithms, assuming that they are used in the protocol itself. Nevertheless, while reading the paper it is most convenient to consider the most practically-efficient naïve algorithm for $n\times k$ and $k\times m$, which runs in $\Theta(nkm)$ time. We remark that the essentially the only non-black-box property of naïve MatMul algorithm the protocol uses is self-reducibility: computing $A\cdot B$ by decomposing it into smaller $(n/r)\times(n/r)$ MatMul of $r\times r$ block matrices (storing intermediate values in the block-wise inner products). This self-reducibility property of MatMul is all the more used in FMM divide-and-conquer algorithms, hence if FMM becomes practical, it should be possible to adapt our protocol to the $n^{\omega}$ baseline as opposed to $n^{3}$.
A formal definition and a framework.
As an independent contribution, to the best of our knowledge, we formulate the first non-trivial definition for proofs of work associated with computational tasks. This definition allows reasoning about different schemes and compare them one to another.
1.2 Related Work
There have been various proposals to completely get rid of the wastefulness of PoW mining by requiring the “waste” of a different resource. Most well known is the notion of “proof of stake” where participants in the network need to lock up a sum of money and thereby receive voting rights proportion to their quantity of holdings (e.g., [BLMR14, GHM+17, KKKZ19] to mentioned just a few). Other resources, such as space [DFKP15, PKF+18, CP19], have been studied and deployed in systems. While these avoid energy waste, they incur waste in other domains, and they all introduce new economic and cryptographic challenges and tradeoffs due to the nature of the other resource being utilized to secure the network.
Trapdoor matrices and algorithmic speedup.
In two recent beautiful works by Vaikuntanathan and Zamir [VZ25] and Braverman and Newman [BN25], the authors discovered a technique for sampling pseudorandom matrices such that whomever has a secret key can multiply them in near-linear time. The suggested constructions rely on standard cryptographic assumptions, i.e., learning parity with noise (LPN). The authors of [VZ25] suggest to use those so-called “trapdoor matrices” in various algorithmic applications that rely on products random matrices in order to obtain a significant speedup; among a series of applications, they emphasize a faster-than-known classification inference algorithm. The authors of [BN25] suggest applications of a similar technique in a context of secure delegation of linear algebra, wherein a (weak) client wishes to compute some public function of its private data. They achieve such protocols for linear public functions (i.e., matrix-matrix multipllication or matrix-vector multiplication) by “masking” the client’s private input with a “trapdoor matrix”.
The above goal is incomparable and orthogonal to proofs of work, and the techniques seem to be unrelated. First, the goal in proofs of work is somewhat reverse: we want hardness to be uniform for everyone, even for those who choose the inputs. Second, in proofs of work there is no possibility for hidden trapdoors as all of the computation is done locally by the party doing the computation. Whether ideas from the above works have any indirect bearing to our protocol is an interesting question.
2 Overview of our PoUW
We first recall the main properties and syntax we would like to have in a PoUW, focusing on the task of matrix multiplication. In this overview, we ignore the economic value of multiplying two matrices, being a central operation in many large scale computations (notably AI training and inference; see Section 7), and focus only on the technical aspects of a PoUW. A PoUW consists of two algorithms, $\mathsf{Solve}$ and $\mathsf{Verify}$, with the following properties, stated informally:
  • $\mathsf{Solve}(\sigma,A,B)$ gets as input a seed $\sigma$ and two $n\times n$ matrices $A,B$ (the instance). The procedure has two outputs: (1) a matrix $C$ and (2) a proof $\pi$.
    Prover efficiency: The running time of $\mathsf{Solve}$ is not much more than just multiplying $A$ and $B$. That is, if $t(n)$ is the time it takes to multiply $A$ and $B$, then $\mathsf{Solve}$ runs in time $t(n)\cdot(1+o(1))$.
    Proof size: The size of $\pi$ is $o(t(n))$.
  • $\mathsf{Verify}(\sigma,\pi)$ gets as input a seed $\sigma$ and a proof $\pi$, and it outputs a bit.
Correctness of a PoUW says that $C=A\cdot B$ (with probability 1) and also $\mathsf{Verify}(\sigma,\pi)=1$ with probability $\epsilon$ (over the randomness of $\mathsf{Solve}$ and over a uniformly random $\sigma$), where $\epsilon>0$ is a parameter. Hardness of a PoUW says that no one can create an alternative $\mathsf{Solve}^{*}$ procedure which is able to run in time $t(n)\cdot(1+o(1))$ and succeed in creating $\pi$’s that are accepted with probability noticeably higher than $\epsilon$. That is, with the same computational budget there is no way to noticeably outperform the honest prover, $\mathsf{Solve}$.
We did not put any restriction on the running time of $\mathsf{Verify}$ in order to emphasize the challenging parts of the notion of a PoUW. Our final scheme will be quite efficient also in communication and verifier complexity; we discuss these later in the overview.
The key idea.
The key idea underlying our PoUW is to encode noise into the input matrices, ensuring that any attempt to “shortcut” the computation would require effort comparable to performing the full computation. At the same time, we design an optimally efficient random self-reduction procedure, incurring negligible computational cost. The latter is obtained by leveraging, not only the task of producing an output, but more importantly, enforcing unpredictability in the transcript of the computation.
A MatMul random-self-reducibility-based PoW.
To gain intuition, we design a proof of work based on multiplying two arbitrary matrices that is not useful. We will discuss later how we make it useful with minimal computational cost. The idea is to ask the prover to compute a “noisy” matrix product, instead of the original one. Specifically, we require computing the following:
$C^{\prime}=(A+E)\cdot(B+F),$ (2.1)
where the matrices $E$ and $F$ are uniformly random $n\times n$ matrices matrices. It is obviously necessary for $E$ and $F$ to be unpredictable by the (malicious) prover, as otherwise it could pre-process the result of (2.1). The latter property can be achieved by deriving $E$ and $F$ by applying a random oracle on $\sigma,A,B$. Since $\sigma$ is not known to the (malicious) prover ahead of time and $\mathcal{O}$ is a random oracle, the output of the random oracle is random and unpredictable. We can now turn this into a proof of (non-useful) work as follows:
  • The prover, on input seed $\sigma$ and two matrices $A,B$, does:
    1. 1.
      Compute $E,F=\mathcal{O}(\sigma,A,B)$.
    2. 2.
      Compute $C^{\prime}=(A+E)\cdot(B+F)$.
    3. 3.
      Compute $z=\mathcal{\mathcal{missing}}O(C^{\prime})$.
    4. 4.
      Output $\pi=(A,B,z)$.
  • The verifier, on input $\sigma$ and $\pi=(A,B,z)$, does:
    1. 1.
      Recompute $z$ from $\sigma,A,B$ to check correctness.
    2. 2.
      Check if $z$ is below a threshold, say $2^{\lambda-\log(1/\epsilon)}$ (assuming $\epsilon$ is the inverse of a power of 2).
Since $E$ and $F$ are uniformly random matrices, the matrix $C^{\prime}$ is uniformly random, no matter what $A$ and $B$ are. Therefore, $z$ will be below the above-mentioned threshold with probability exactly $\epsilon$, as needed. Suppose that $t(n)$ is the time spent by the prover on matrix multiplication in step 2. Note that $t(n)=O(n^{3})$ if we use classical matrix multiplication, and $t(n)=n^{\omega}$ if we use fast matrix multiplication, where the current record is $\omega<2.371339$ [ADV+24]. Besides step 2, all operations are linear time in the input size and are thus negligible. Overall, the prover’s runtime in the above system is $t(n)\cdot(1+o(1))$.
For the above system to be a PoW, we need to argue that there is no way to perform noticeably better than the honest prover. In our system, the verifier checks consistency of $z$ and so the only way to perform better is if an attacker could somehow come up with $A$ and $B$ for which it can compute $z$ faster than doing a generic matrix multiplication algorithm. Suppose otherwise, then we could use such a prover to compute the product of two random matrices in time $o(t(n))$ — a major breakthrough in algorithms. Assuming no such speedup is possible for random matrices, we conclude that a malicious prover must run in time at least $t(n)$, as much as the honest prover one up to additive lower-order terms.
Obviously the above PoW is not useful since the prover does not output the product of $A$ and $B$. Instead, it can only output a noisy and seemingly useless value $C^{\prime}=(A+E)\cdot(B+F)$. Naively, one can add another step to the prover’s algorithm in which it computes the “noise” term $A\cdot F+E\cdot(B+F)$ and subtracts it from $C^{\prime}$ to output $C=A\cdot B$. Computing this term, however, requires two additional generic matrix products (plus lower-order tasks). Thus, the honest prover would need to perform $3$ matrix products, resulting in a prover whose running time is $t(n)\cdot(3+o(1))$, while a malicious prover still needs to do only a single matrix product for the same chance of convincing the verifier. Hence, the above scheme is not a PoUW.
Usefulness via transcript unpredictability.
The issue with the above PoW not being useful was that “peeling off” the noise was too computationally expensive. Let us explore a different method of adding noise for which peeling it off is going to be relatively easy. Instead of sampling the noise matrices $E$ and $F$ as completely uniform, we sample them as low rank with parameter $r\ll n$. It is straightforward to do this by sampling two matrices $E_{L},F_{L}$ of size $n\times r$ and two matrices $E_{R},F_{R}$ of size $r\times n$, and then computing $E=E_{L}\cdot E_{R}$ and $F=F_{L}\cdot F_{R}$. As a result, $E$ and $F$ are two $n\times n$ random rank $r$ matrices. While $r$ is a generic parameter, for this overview, imagine that $r=n^{0.3}$ for convenience.
Notice that “peeling off” the noise in $C^{\prime}$ is now computationally easy. Indeed, computing $A\cdot F+E\cdot(B+F)$ (cf. Eq. (2.1)) can be done by two matrix product where in each of them, one of the matrices is low rank, allowing us to compute the product in time $O(n^{2}\cdot r)$. This term is asymptotically negligible compared to the cost of “standard” matrix multiplication, i.e., $O(n^{3})$ using classical matrix multiplication or $n^{\omega}$ using fast matrix multiplication.
Once we have managed to get the correct output of the matrix product, the question is whether the new scheme is even a proof of work, or perhaps because $E$ and $F$ are “less random” an attacker can now convince the verifier with noticeably higher probability (and with the same computational effort) than the honest prover. The latter is unfortunately indeed the case: a malicious prover can choose easy to multiply matrices $A$ and $B$, say the all 0 matrices, and then its computational task boils down to merely multiplying $E$ and $F$. Being low-rank matrices, the adversary can compute $E\cdot F$ and then feed the output to the random oracle in time roughly $O(n^{2}\cdot r)$ which is much faster than the honest prover, requiring time $O(n^{3})$ (or $n^{\omega}$ using fast matrix multiplication). It may seem as if we have not made any progress, but as we explain next, this is not the case.
As we concluded above, since $E$ and $F$ are low rank, we cannot utilize the output of the computation as a source of hardness. To this end, we make the following novel observation: instead of using the output of the computation as the proof of work, we utilize the transcript of the computation. This essentially ties the hands of the attacker and forces it to follow a prescribed algorithm, as we explain next.
Consider the algorithm that multiplies $A^{\prime}=(A+E)$ and $B^{\prime}=(B+F)$ using a block-variant of the classical algorithm. Specifically, we split $A^{\prime}$ and $B^{\prime}$ into $r\times r$ blocks and initialize an all 0s matrix ${C^{\prime}}^{(0)}$. We then iteratively compute the product of $A$ and $B$ via the following formula
$C^{\prime(\ell)}_{i,j}:=C^{\prime(\ell-1)}_{i,j}+A_{i,\ell}\cdot B_{\ell,j},$
where we treat the $(i,j)$-th index of each matrix as an $r\times r$ submatrix (or block). Namely, $i,j$, and $\ell$, all range in $1,\ldots,n/r$ (and we assume $r\mid n$ for simplicity). Obviously, $C^{\prime}$ consists of the blocks $\{{C^{\prime}}^{(n/r)}_{i,j}\}_{i,j\in[n/r]}$.
This sequence of $(n/r)^{3}$ matrices $\{{C^{\prime}}^{(n/r)}_{i,j}\}_{i,j\in[n/r]}$ is referred to as the transcript of the matrix product. Our proof of work is then obtained by applying a random oracle on the transcript, instead of on the output. Specifically, we suggest the following PoUW:
  • The prover, on input seed $\sigma$ and two matrices $A,B$, does:
    1. 1.
      Compute $E_{L},E_{R},F_{L},F_{R}=\mathcal{O}(\sigma,A,B)$ such that $E=E_{L}\cdot E_{R}$ and $F=F_{L}\cdot F_{R}$ and they are of dimension $n\times n$.
    2. 2.
      Compute the transcript of $C^{\prime}=(A+E)\cdot(B+F)$, denoted $\mathsf{tr}=\{{C^{\prime}}^{(n/r)}_{i,j}\}_{i,j\in[n/r]}$.
    3. 3.
      Compute $z=\mathcal{\mathcal{missing}}O(\mathsf{tr})$.
    4. 4.
      Compute $C=C^{\prime}-(A\cdot F+E\cdot(B+F))$.
    5. 5.
      Output $C$ and $\pi=(A,B,z)$.
  • The verifier, on input $\sigma$ and $\pi=(A,B,z)$, does:
    1. 1.
      Recompute $z$ from $\sigma,A,B$ to check correctness.
    2. 2.
      Check if $z$ is below a threshold, say $2^{\lambda-\log(1/\epsilon)}$ (assuming $\epsilon$ is the inverse of a power of 2).
As we have already explained, the above scheme satisfies correctness in the sense that a verifier will accept an honestly generated proof with probability $\epsilon$ and also $C$ is equal to $A\cdot B$. For efficiency, notice that steps 1 and 4 can be implemented in time $O(n^{2}\cdot r)$ and step 3 requires hashing a string of length $(n/r)^{3}$ which takes time $o(n)$. Step 2 requires computing the whole transcript of the matrix multiplication (which includes the output, $C^{\prime}$). One way to compute it is directly running the classical matrix multiplication algorithm (in time $O(n^{3})$). However, there is a more efficient way using fast matrix multiplication techniques. Indeed, every intermediate $r\times r$ block can be computed via a fast rectangular matrix multiplication that can be done non-trivially in some regimes of parameters. For instance, one can multiply a $n\times n^{0.3}$ matrix with a $n^{0.3}\times n$ matrix in time $n^{2+o(1)}$ [VXXZ24]. Thus, in some regimes one can implement step 2 (and so step 3) in time $O(n^{3}/r)$. Depending on $r$, let us denote the current state of the art runtime of the above algorithm for step 2 by $t_{r}(n)$. Since the latter is the dominant term, we get that the total runtime of the honest prover is $t_{r}(n)\cdot(1+o(1))$.
For hardness, we conjecture that the algorithm we present for the honest prover is essentially optimal (up to low order optimizations) for an attacker whose goal is to cause the verifier to accept. Indeed, an attacker can choose $A$ and $B$ and completely avoid running step 4, but we conjecture that there is no noticeable speedup possible in the computation of step 2.
To this end, we formalize an assumption saying roughly that computing the transcript of a product of two random rank $r$ matrices requires time $t_{r}(n)$. With this conjecture, we conclude hardness of our scheme. We now give reasons to believe this conjecture, namely, why computing the transcript requires essentially to compute every $r\times r$ intermediate matrix independently. The main point is that every intermediate ${C^{\prime}}^{(\ell)}_{i,j}$ essentially depends on the product of two $r\times r$ blocks from $A^{\prime}=A+E$ and $B^{\prime}=B+F$, denoted $A^{\prime}_{i,\ell}$ and $B^{\prime}_{\ell,j}$. Since $E$ and $F$ were chosen to be uniformly random low rank matrices, we can infer that both $A^{\prime}_{i,\ell}$ and $B^{\prime}_{\ell,j}$ are completely uniform, marginally. Thus, computing any single intermediate does require doing a single $r\times r$ random matrix multiplication, which as we conjectured cannot be sped up (from $O(r^{3})$ directly or from $O(r^{\omega})$ with fast matrix multiplication). Obviously the above argument does not work for computing all of the intermediates because they are correlated (remember that $E$ and $F$ are only rank $r$). However, recovering the correlations requires essentially solving a linear system of equations which, we conjecture, is as difficult as multiplying matrices. Thus, we believe our conjecture holds, unless significant algorithmic breakthrough is obtained.
Remark 2.1 (Memory complexity).
As described the memory consumption of the PoUW is quite significant due to the need to store the whole transcript $\mathsf{tr}$. However, there are ways to improve this. For example, instead of hashing the whole transcript, one can hash each of the $(n/r)^{3}$ intermediate matrices separately, and use each one of them as a potential proof for solving the puzzle. The latter, in turn, causes the number of attempts one gets in solving the puzzle scale proportionally with the size of the matrices they multiply. This is a useful feature if the system needs to support varying matrix sizes. Additionally, note that in terms of hardness, a similar argument to the above applies in this case: each intermediate matrix is the result of the product of two independent, each marginally uniformly random matrices, causing the output to be “computationally” unpredictable (in a fine-grained sense).
3 Open Problems and Future Directions
The core idea in this paper (in particular, the PoUW Scheme in Algorithm 6.5) is performing PoW on intermediate computations as opposed to the output of the task $A\cdot B$. Once one subscribes to this idea and to a stronger security assumption (direct-product of random but correlated subproblems), it is natural to ask how general is our approach:
Problem 3.1.
What other computational tasks $f(x)$ beyond MatMul admit an efficient ($1+o(1)$ multiplicative overhead) PoUW scheme? Can we characterize the properties of functions $f(x)$ for which an efficient PoUW scheme exists?
Some natural and interesting candidates are Linear Equations, Graph search and routing problems, and database problems like Pattern-Matching and string problems (e.g., DNA sequencing). Some of these functions lack a natural decomposable structure into sub-problems $f(x)=\bigotimes g(x_{i})$ like MatMul, so it is less clear how to use our notion of random self-reducibility on subproblems.
Alternative Noise schemes for MatMul PoUW
Our PoUW scheme in Algorithm 6.5 is merely one simple instance of an efficient (sub-cubic overhead) random self-reduction for MatMul, but there is in principle “nothing holy" about using low-rank noise – Indeed, all we really need the noise structure to satisfy is that it is: (i) easy to add and remove ($o(n^{3})$), (ii) induces unpredictable marginals (random) on tiles, no matter how $A,B$ are chosen. A natural question is whether there are even more efficient (and secure) schemes:
Problem 3.2.
Is there a better PoUW scheme for MatMul than Algorithm 6.5, either in terms of computational overhead ($O(n^{2}r)$) or the security assumption (ideally both)?
In Appendinx A we present an alternative candidate PoUW scheme with self-canceling noise, i.e., which preserves the output $AB$ of the original (useful) problem, without any need for "denoising" altogether. This scheme has $O(n^{2}\log n)$ computational overhead (at least for certain matrix dimensions) instead of $O(n^{2}r)$ of Algorithm 6.5, but only applies to high-rank matrices (typically satisfied by useful real-world matrices), and relies on a more subtle security assumption. Nevertheless, it demonstrates the flexibility of random self-reductions for MatMul, and the generality of our PoUW framework.
4 Preliminaries
Notation.
For an integer $n\in\mathbb{N}$, we denote $[n]=\{1,\ldots,n\}$. For a set $D$, $x\leftarrow D$ denotes sampling an element from $D$ uniformly at random. For a distribution $\mathcal{D}$ over a set $D$, $x\leftarrow\mathcal{D}$ denotes sampling an element of $D$ according to $\mathcal{D}$. By $U_{m}$ we denote the uniform distribution over $\{0,1\}^{m}$.
Random oracle model and complexity.
Throughout this paper we assume the random oracle model (ROM). Proof of work protocols are easiest to define and prove secure in this model because it allows counting a query to the RO as a single operation/time unit. Nevertheless, everything can be heuristically applied to the standard model, by replacing the random oracle with a cryptographic hash function (say SHA-256 or BLAKE3).
In addition to supporting random oracle queries, we allow parties (as usual) to perform arithmetic operations. We assume an underlying field $\mathbb{F}_{q}$ that supports addition and multiplication. We count complexity of algorithm in our model by separately accounting for the number of random oracle queries they make and the number of field operations they do.
5 Definitional Framework for Proofs of Useful Work
A proof of useful work (PoUW) is a method for a prover to convince a verifier that it performed sufficient computation to solve a particular computational task. Importantly, the task the prover chooses to solve is arbitrary, perhaps chosen by the prover itself, and does not come from any pre-specified distribution. The task may even remain private, depending on the application. Soundness requires that any (malicious) prover running sufficiently faster than the honest prover will not be able to fool the verifier to accept with high probability (even if they completely control the task they solve). More precisely, the ratio between the amount of computational effort and the probability of convincing a verifier is very close to the honest party’s ratio). In other words, a proof of useful work has to satisfy the following two properties: aim to achieve the following two desiderata simultaneously:
  1. 1.
    Usefulness: Acceptable proofs can be generated by solving “useful” computational problems on arbitrary inputs.
  2. 2.
    Efficiency: The (honest) prover needs to be almost as efficient as only solving the task (without proving).
  3. 3.
    Hardness: Generating acceptable proofs is guaranteed to necessitate actual work, similar to the amount of work required honestly.
Classical proofs of work [DN92] only require the last property to hold, and have no associated useful computation. Usually, one can leverage classical proofs of work to generate random instances of “useful” problems by sampling them from the proof of work challenge. In this way, for appropriate tasks (where average case complexity is well understood), a prover is guaranteed to spend a certain amount of work in producing proofs. However, while the algorithm being executed for generating a proof of work is related to a concrete computational task, the instance itself is random and detached from any fixed instance that someone may actually want to solve. Thus, such systems are not considered useful according to the above.
We formalize a proof of useful work as a protocol between a prover $P$ and a verifier $V$. At a high level, $P$ has an instance of a task that it needs to solve. Depending on the application, the instance could very well be already in $P$’s possession, or alternatively, may be given to $P$ from an external entity. Then, $P$ and $V$ engage in a protocol where eventually $P$ outputs a solution to the task and $V$ gets convinced that $P$ spent sufficient computational effort.
It is very common to see definitions of proofs of work as non-interactive algorithms rather than interactive protocols. We may jump back and forth between the two methods, depending on the context. We remark that all of our protocols are public-coin and therefore we can use the Fiat-Shamir heuristic [FS86] to remove the interaction and prove security in the Random Oracle (RO) model . Thus, in what follows we assume the RO model, meaning that all parties have query access to a public random function.
Since our proofs of work are associated with a useful computation, we assume a function $f\colon\{0,1\}^{n}\to\{0,1\}^{n}$, reflecting the computation being done during the proof of work. Above, we assume that $f$ has the same domain and range for simplicity, and all of our definitions extend to the case where the domain and range differ. Further, all of our definitions extend to the case where $f$ could have several “legal” outputs per input, namely, where $f$ is a relation. We assume that $f$ is a function for simplicity. We assume that there is a canonical algorithm Alg for solving $f$, namely, the algorithm gets as input $\mathsf{x}\in\{0,1\}^{n}$ and outputs a solution $\mathsf{y}\in\{0,1\}^{n}$ such that $f(\mathsf{x})=\mathsf{y}$. Denote $t(n,\mathsf{x})$ the complexity of Alg in solving $\mathsf{x}\in\{0,1\}^{n}$. We denote $t(n)=\max_{\mathsf{x}}f(n,\mathsf{x})$ the worst-case complexity of Alg over all inputs of size $n$.
Throughout we assume that $\lambda\in\mathbb{N}$ is a security parameter, and assume that $\{0,1\}^{\lambda}$ is the domain and range of the random oracle. One should think of $n\gg\lambda$ where the relationship between the two is via some polynomial. Syntactically, the definition of a PoUW involves two oracle-aided algorithms:
  1. 1.
    $\mathsf{Solve}(\sigma,\mathsf{x})$: is a probabilistic algorithm that takes as input a seed $\sigma\in\{0,1\}^{\lambda}$ and an instance $\mathsf{x}\in\{0,1\}^{n}$, and outputs a solution $\mathsf{y}\in\{0,1\}^{n}$ and a proof $\pi\in\{0,1\}^{*}$.
  2. 2.
    $\mathsf{Verify}(\sigma,\pi)$: is a deterministic algorithm that takes a seed $\sigma\in\{0,1\}^{\lambda}$ and a proof $\pi\in\{0,1\}^{*}$ as input, and outputs 0 (for failure) or 1 (for success).
The role of $\sigma$ is to guarantee that the computational work spent in order to “solve” a task is “fresh,” i.e., executed when $\sigma$ became publicly known to avoid preprocessing. One can further assume that there is a procedure $\mathsf{Keygen}$ that outputs $\sigma$, and also possibly a $\mathsf{GenPP}$ procedure that outputs public parameters $\mathsf{pp}$ that all algorithms get as input. We simplify and avoid formally stating these for succinctness and clarity.
We capture the efficiency, completeness, and soundness properties in the next definition.
Definition 5.1 (Proof of useful work).
A PoUW for $f$ consists of two algorithms $\mathsf{PoUW}=(\mathsf{Solve},\mathsf{Verify})$. These algorithms must satisfy the properties $(\alpha,\beta)$-efficiency, $\epsilon$-completeness, hardness for functions $\alpha,\beta,\epsilon$, as described next:
$(\alpha,\beta)$-efficiency: For any $\sigma\in\{0,1\}^{\lambda}$ and any $\mathsf{x}\in\{0,1\}^{n}$:
  • $\mathsf{Solve}(\sigma,\mathsf{x})$ runs in time $\alpha(t(n))$.
  • $\mathsf{Verify}(\sigma,\pi)$ runs in time $\beta(t(n))$ for any possible output $\pi$ of $\mathsf{Solve}$.
$\epsilon$-completeness: For every $\lambda,n\in\mathbb{N}$, every $\sigma\in\{0,1\}^{\lambda}$, and every $\mathsf{x}\in\{0,1\}^{n}$, it holds that
$\Pr[f(\mathsf{x})=\mathsf{y}]=1\;\text{and}\;\Pr[\mathsf{Verify}(\sigma,\pi)=1]\geq\epsilon(n),$
where the probabilities are over the randomness of $\mathsf{y},\pi\leftarrow\mathsf{Solve}(\mathsf{x},\sigma)$ and over the random oracle.
Hardness: For any $\delta>0$ and any algorithm $\mathsf{Solve}^{\star}$ with unbounded polynomial-time preprocessing and whose online running time is $(\alpha(t(n)))^{1-\delta}$, it holds that
$\Pr\left[\mathsf{Verify}(\sigma,\pi)=1\right]\leq\epsilon(n),$
where the probability is over the randomness of $\sigma\leftarrow\{0,1\}^{\lambda}$ sampled independently uniformly at random, computing $\pi\leftarrow\mathsf{Solve}^{\star}(\sigma)$, and over the random oracle.
A PoUW scheme is said to be optimal if $\mathsf{Solve}$ introduces minimal overhead over directly solving $\mathsf{x}$, and if no adversary can obtain noticeable advantage over the honest prover. We also require non-trivial efficiency on the verifier. We make this explicit in the next definition.
Definition 5.2 (Optimal PoUW).
A PoUW for $f$, denoted $(\mathsf{Solve},\mathsf{Verify})$ is said to be optimal if
  • $\alpha(t(n))=t(n)\cdot(1+o(1))$ and $\beta(t(n))=t(n)\cdot(1-\omega(1))$
  • $\epsilon$ is any constant between 0 and 1.
We conclude this section with several remarks about the definition.
Remark 5.3 (Tunable difficulty).
In applications it is often necessary to be able to tune the hardness of the problem to solve. For example, in Bitcoin, the more energy the prover invests, the higher are the chances of convincing the verifier; moreover, once in a while the overall difficulty of mining becomes higher, requiring more energy to mine a block.
This can be achieved in two ways. One way is to have $\epsilon$ fixed, say $0.1$, and vary the size of the tasks $\mathsf{x}$ the prover need to solve. Another way is to fix $n$ and tune $\epsilon$, requiring to solve “more instances” to achieve the same probability of convincing the verifier.
In our final protocol, both $n$ and $\epsilon$ are tunable, and so one can either vary $n$ (choose smaller or larger instances to solver) or vary $\epsilon$ (require less or more probability of success) to tune the difficulty of the prover. In the above definition we treat $\epsilon$ as a fixed function only for simplicity.
Remark 5.4 (One vs. two-sided correctness).
We require that the (honest) prover outputs the correct $\mathsf{y}$ for each $\mathsf{x}$ it wants to solve (with probability 1). We do so for simplicity and because our construction will satisfy this strong requirement. However, it is plausible to consider a weaker definition where a correct solution would be outputted only with some reasonably high probability.
Remark 5.5 (Prover’s efficiency).
We parameterize the efficiency of $\mathsf{Solve}$ using a general function $\alpha$. However, for practical purposes, we believe that the overhead of $\mathsf{Solve}$ should be as small as possible. Indeed, with constant multiplicative overhead, there is already a PoUW where solving the task and proving to the verifier that work has been done are done separately and independently (see Section 5.1 for details). Thus, the interesting regime is where $\alpha(t(n))\leq c\cdot t(n)$ where $c\ll 2$. In our protocol, we will achieve the best one can hope for $\alpha(t(n))\leq t(n)\cdot(1+o(1))$.
Remark 5.6 (On hardness).
Hardness considers a scenario where an attacker gets a fresh challenge of the PoUW. We provide the prover with a complexity budget sufficient to run the honest prover and ask it to perform significantly better, i.e., have much better than $\epsilon$ probability of causing the verifier to accept. In words, this should be interpreted as saying that the honest provers strategy is optimal, up to small terms.
One can introduce stronger notions of security, considering the success probability of a malicious prover when working on multiple instances in parallel (i.e., whether amortization exists), or when the adversary receives challenges in a streaming fashion and it only needs to show slight advantage over the naive process. The latter property is essentially what is needed for Bitcoin-style blockchains, and we formalize the security definition in Appendix B We conjecture that our PoW mechanism will satisfy all properties satisfied by the PoW of Bitcoin, and focus in the main body on the simpler definition to simplify presentation and convey the main new ideas.
Remark 5.7 (Additional constraints).
One could impose additional constraints on the efficiency of the protocol, e.g., that the proof $\pi$ sent from $P$ to $V$ is very short, or that the verifier’s running time is sufficiently small. We will discuss these later in the paper.
5.1 A (Trivial) PoUW for Any Function
We show that a PoUW for any function $f$ with an associated Alg running in time $t(n)$ on instances of size $n$ exists, albeit this scheme is not optimal. Specifically, there is a gap between the complexity of $P$ and of the associated algorithm, and also there is a $P^{\star}$ that has significantly less complexity than $P$. At a very high level, the scheme is doing a computation plus proof of work, independently. More details follow.
  • $\mathsf{Solve}$: on input $\sigma$ and $\mathsf{x}$, the prover solves the task by running Alg on $\mathsf{x}$ to get $\mathsf{y}$. Also, independently, it does a proof of work to convince the verifier that it spent enough time. Specifically, it chooses $T=c\cdot t(n)$ nonces $\omega_{1},\ldots,\omega_{T}$ for the random oracle, and computes $z_{i}=\mathcal{O}(\sigma,\omega_{i})$. It then sets $\pi$ to be the $\omega_{i}$ for which $z_{i}$ starts with the most 00’s.
    The prover’s running time is thus $\alpha(t(n))=(1+c)\cdot t(n)$.
  • $\mathsf{Verify}$: on input $\sigma$ and $\omega$, the verifier outputs $1$ if and only if $\mathcal{O}(\sigma,\omega)$ has $\log(t(n))$ leading 00’s.
    The verifier’s running time is thus $\beta(t(n))=\tilde{O}(1)$.
Claim 5.8 (A PoUW for $f$).
The above PoUW protocol satisfies $(1-\epsilon)$-completeness and hardness (see Definition 5.1) with $\epsilon=e^{-c}$.
Proof.
Since $\mathcal{O}$ is a random oracle, its output on previously unqueried points is uniformly distributed in $\{0,1\}^{\lambda}$. Furthermore, the probability of success for a single query is $p=1/t(n)$. The probability that all $T$ attempts fail, is $(1-p)^{c\cdot t(n)}\leq e^{-c}$. Thus, $(1-\epsilon)$-correctness of proof holds. Hardness follows directly, again, from the random oracle property. A method for “guessing” a nonce that leads to an output that has $\log T$ leading 0s amounts to “guessing” the output of a random oracle without even making the appropriate query. ∎
The protocol is not an optimal PoUW. Indeed, either $P$ has linear overhead over the complexity of Alg, or an adversarial $P^{\star}$ can gain linear advantage in running time over $P$. Details follow. Consider the following malicious prover $P^{\star}$: $P^{\star}$ does not multiply any two matrices, and instead only finds the good nonce out of $T$ arbitrary ones. Clearly, this $P^{\star}$ has exactly the same probability of convincing $V$ as $P$, but it is more efficient. Concretely, in this protocol protocol $t_{P^{\star}}(n)=c\cdot t(n)$ which is a factor $(1+1/c)$ faster than $t_{P}(n)$. It is possible to make the advantage of $P^{\star}$ sub-linear by setting $c=\omega(1)$, but this takes the other issue to the extreme: the honest prover $P$’s complexity is a multiplicative factor $(1+c)$ larger than the complexity of Alg.
6 A PoUW for Matrix Multiplication
In this section we present a PoUW protocol for the matrix multiplication task. In this context, the task consists of two $n\times n$ matrices $A$ and $B$ with entries integers in $\mathbb{F}_{q}$ (we assume field operations), and the solution is an $n\times n$ matrix $C$ being the product of $A$ and $B$ with entries also being in $\mathbb{F}_{q}$. [^4] The protocol can be extended to rectangular matrix multiplication, but again we favor simplicity of presentation.
This section is structured as follows. We first present the canonical matrix multiplication algorithm that we assume (referred to as MatMul). Our protocols are roughly based on the random self reducibility property of matrix multiplication. Our first protocol is rather straightforward, and while being better than the generic one in Section 5.1, it is still not an optimal PoUW. Our second protocol combines the idea of random self reducibility of matrix multiplication together with an additional structural property of the algorithm we us. This protocol, as we show, is an optimal PoUW under a new algorithmic conjecture.
6.1 The Canonical MatMul Algorithm
For concreteness, we describe the textbook algorithm for matrix multiplication, denoted MatMul. In this algorithm, the $(i,j)$ entry of the output is computed by performing an inner product between the $i$-th row in the left matrix and the $j$-th column in the right matrix. We generalize this algorithm to operate on $r\times r$-size blocks (or so-called “tiles”), instead of on single entries. See Algorithm 6.1 for the pseudocode.
{mdframed}
[ backgroundcolor=codeBackground, linecolor=black, linewidth=0.5pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]Canonical Matrix Multiplication Algorithm
[…]
  1. 1.
    Initialize an all-00 matrix $C^{(0)}\in[-q,q]^{n\times m}$.
  2. 2.
    Partition $A$, $B$, and $C$ into blocks of size $r\times r$. The $(i,j)$th block of a matrix $X\in\{A,B,C\}$ is denoted $X_{i,j}$.
  3. 3.
    For each $i\in[n/r]$, $j\in[m/r]$, and $\ell\in[k/r]$, compute
    $C^{(\ell)}_{i,j}:=C^{(\ell-1)}_{i,j}+A_{i,\ell}\cdot B_{\ell,j},$
    where $A_{i,\ell}\cdot B_{\ell,j}$ can be computed via any classical matrix multiplication algorithm.
  4. 4.
    Output $C^{(n)}$.
Algorithm 6.1
The complexity of $\textsc{MatMul}_{r}$, when multiplying an $n\cdot k$ matrix with a $k\times m$ matrix, is governed by the computation of the $nkm/r^{3}$ “intermediate’ values. When we implement the computation of the intermediate values naively, we obtain an algorithm with complexity $t_{\textsc{MatMul}}(n,k,m)=O(n\cdot k\cdot m)$. To simplify notation, when we multiply two $n\times n$ matrices, i.e., $n=k=m$, we write $t_{\textsc{MatMul}}(n)=t_{\textsc{MatMul}}(n,k,m)$. We assume throughout that $r$ divides $n,k,m$.
$\textsc{MatMul}_{r}$ transcript.
The transcript of a $\textsc{MatMul}_{r}$ (Algorithm 6.1) execution consists of all $nmk/r^{3}$ intermediate $r\times r$ matrices $\{C^{(k)}_{i,j}\}_{i\in[n/r],j\in[m/r],\ell\in[k/r]}$ during the computation of $A\cdot B$ for $A\in{\mathbb{F}}_{q}^{n\times k}$ and $B\in{\mathbb{F}}_{q}^{k\times m}$.
Definition 6.1 (Transcript).
The transcript of $\textsc{MatMul}_{r}(A,B)$ consists of
$\mathsf{Tr}(\textsc{MatMul}_{r},A,B)=\left\{C^{(k)}_{i,j}\right\}_{i\in[n/r],j\in[m/r],\ell\in[k/r]}.$
Evidently, in general, there are $nmk/r^{3}$ intermediate matrices, each of which is of size $r^{2}$. We give some examples, assuming that $n=m=k$ for simplicity. If $r=n$, there is exactly 1 intermediate state and it is the output of the computation. Alternatively, if $r=n/2$, there are 8 intermediate states, two states per one of the 4 sub-rectangles of $C$. Lastly, if $r=1$, then there are $n^{3}$ intermediate matrices, each being a single value in ${\mathbb{F}}_{q}$.
Algorithms for computing the intermediates.
As mentioned, one can compute all of the $nkm/r^{3}$ intermediate matrices using a direct matrix multiplication algorithm, requiring $O(r^{3})$ work per intermediate matrix, resulting with overall complexity $O(nmk)$. However, using fast matrix multiplication techniques one can do better.
The first idea is to use fast square matrix multiplication algorithms. Specifically, every intermediate value is obtained by adding to a previously calculated intermediate value the product of two $r\times r$ matrices. Since this can be done in time $r^{\omega}$ where $\omega$ is the exponent of matrix multiplication (currently standing at $\omega=2.371552$ [VXXZ24]), we obtain an algorithm with overall complexity $O(nmk/r^{3-\omega})$, beating the naive one.
A somewhat more efficient approach in many cases is to utilize fast rectangular matrix multiplication algorithms. Specifically, consider the matrix induced by the values $C_{i,j}^{\ell}$ for a fixed $\ell$. Let us focus on $\ell=1$ for concreteness. This matrix is, in fact, the product of the left-most column-strip of $A$ (when we view the matrix as having $r\times r$ atomic blocks as elements) and the top-most row-strip of $B$ (essentially computing the outer product of these vectors). We shall denote $\omega_{m}$ the exponent of the best matrix multiplication algorithm for multiplying $n\times m$ and $m\times n$ matrices for $m<n$. We get an algorithm that runs in time $(n/r)\cdot n^{\omega_{r}}=n^{\omega_{r}+1}/r$, a term that could be $o(n^{3})$. As a concrete setting, the current record says that computing an $n\times n^{k}$ times an $n^{k}\times n$ matrix for all $k\leq 0.321334$ can be done in time $n^{2+o(1)}$ [Gal24, VXXZ24]. That is $\omega_{k}=2$ for all $k\leq 0.321334$. For $r=n^{0.3}$, this results with an algorithm that works in time $n^{2.7+o(1)}$ for computing all of the intermediates.
To conclude, there are several ways to compute the intermediate values, some of which are asymptotically more efficient than the direct method. However, in practice the most efficient algorithm for reasonable values of $n$ is still the direct one because the other ones have either large hidden constants or require significant resources other than compute (e.g., much more memory accesses or an inherent sequentiality).
6.2 A Transcript Unpredictable Encoding Scheme
[…]
  • $A^{\prime},B^{\prime}\Leftarrow\textsc{Encode}_{r}(\sigma,A,B)$: On input a seed $\sigma$ and two matrices $A$ and $B$, it outputs two matrices $A^{\prime}$ and $B^{\prime}$.
  • $C\Leftarrow\textsc{Decode}_{r}(\sigma,C^{\prime})$: On input a seed $\sigma$ and a matrix $C^{\prime}$, the procedure outputs a matrix $C$.
[…]
Definition 6.2 (Transcript-unpredictable encoding).
[…]
Completeness: For every two matrices $A$ and $B$, and auxiliary input $r$, it holds that
$\Pr_{\sigma}[\textsc{Decode}_{r}(\sigma,A^{\prime}\cdot B^{\prime})=A\cdot B\;\colon\;A^{\prime},B^{\prime}=\textsc{Encode}_{r}(\sigma,A,B)]=1.$
$\epsilon$-transcript unpredictability: For every two $n\times n$ matrices $A,B$, and every $r$, every iterative numerical algorithm $\mathcal{A}$ that runs in time $o(t_{\textsc{MatMul}}(n))$ cannot compute the transcript of the product of the noisy matrices $A^{\prime},B^{\prime}$, except with probability $\epsilon$. That is,
$\Pr\left[\begin{aligned} &\forall i,j,k\in[n/r]\colon\\ &\quad C^{(k)}_{i,j}=Z^{(k)}_{i,j}\end{aligned}\,\middle|\,\begin{aligned} &A^{\prime},B^{\prime}=\textsc{Encode}_{r}(\sigma,A,B)\\ &\{C^{(k)}_{i,j}\}_{i,j,k\in[n/r]}=\mathsf{Tr}(\textsc{MatMul}_{r},A^{\prime},B^{\prime})\\ &\mathcal{\{}Z^{(k)}_{i,j}\}_{(i,j,k)\in[n/r]}=\mathcal{A}(A^{\prime},B^{\prime})\\ \end{aligned}\right]\leq\epsilon,$
where the probability is over the choice of $\sigma\leftarrow\{0,1\}^{\lambda}$ and the internal randomness of $\mathcal{A}$.
6.3 From Transcript Unpredictability to a PoUW
An implication of transcript unpredictability is that computing the whole transcript requires (essentially) to perform the direct computation in time $t_{\textsc{MatMul}}(n)$. This directly gives us a PoUW, as described next.
{mdframed}
[ backgroundcolor=codeBackground, linecolor=black, linewidth=0.5pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]A PoUW for Matrix Multiplication with Parameter $r$
[…]
  1. 1.
    Compute $A^{\prime},B^{\prime}=\textsc{Encode}_{r}(\sigma,A,B)$.
  2. 2.
    Computes $C^{\prime}=\textsc{MatMul}_{r}(A^{\prime},B^{\prime})$.
    Denote $z=\{C^{(k)}_{i,j}\}_{i,j,k\in[n/r]}$ the intermediate $r\times r$ matrices.
  3. 3.
    Compute $C:=\textsc{Decode}_{r}(\sigma,C^{\prime})$.
  4. 4.
    Output $(C,\pi)$ where $\pi=(A,B,z)$.
$\mathsf{Verify}(\sigma,\pi)$:
  1. 1.
    Parse $\pi=(A,B,z)$.
  2. 2.
    Recomputes from $\sigma,A,B$ the correct value of $z$, and output $1$ if and only if $z$ is correct.
Algorithm 6.2
Correctness.
It is straightforward to verify correctness. Specifically, the output of $P$ is always $C=A\cdot B$ directly by the correctness of the algorithms Encode and Decode. Also, the verifier accepts because it essentially repeats the computation of the honest $P$, necessarily ending up with the same $z$.
Hardness.
[…]
6.4 A First Instantiation
Our first instantiation of the Encode and Decode operations is quite standard and relies on the classical random self-reducibility of matrix multiplication. In this instantiation we use $r=1$.
{mdframed}
[ backgroundcolor=codeBackground, linecolor=black, linewidth=0.5pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]First Implementation of Encode and Decode
$\textsc{Encode}_{r}(\sigma,A,B)$: $\quad\sslash\quad A,B\in{\mathbb{F}}_{q}^{n\times n}$
  1. 1.
    Parse $\sigma=E,F$, where $E,F\in{\mathbb{F}}_{q}^{n\times n}$.
  2. 2.
    Output $A^{\prime}=A+E$ and $B^{\prime}=B+F$.
$\textsc{Decode}_{r}(\sigma,C^{\prime})$: $\quad\sslash\quad C^{\prime}\in{\mathbb{F}}_{q}^{n\times n}$
  1. 1.
    Parse $\sigma=E,F$, where $E,F\in{\mathbb{F}}_{q}^{n\times n}$.
  2. 2.
    Compute $C^{\prime\prime}=A\cdot F+E\cdot(B+F)$ (using two invocations of $\textsc{MatMul}_{r}$ and a matrix addition).
  3. 3.
    Output $C=C^{\prime}-C^{\prime\prime}$.
Algorithm 6.3
It is obvious that completeness holds as
$C^{\prime}-C^{\prime\prime}=(A+E)\cdot(B+F)-A\cdot F+E\cdot(B+F)=A\cdot B.$
(notice that the $E$ and $F$ matrices are consistent between the two procedures). We now discuss transcript unpredictability. Since $r=1$, the transcript of the computation consists of only the final output $C^{\prime}$. For trace unpredictability, we essentially assume that matrix multiplication for random matrices requires $\Omega(n^{\omega})$ time where $\omega$ is the matrix multiplication exponent. We formalize this next.
Assumption 6.3.
There is no algorithm to compute the product of two random $n\times n$ matrices in time $o(n^{\omega})$.
Efficiency of PoUW.
The obvious downside is that a malicious worker can skip the Decode operation altogether and thereby save 2 matrix multiplications. In other words, the honest worker does at least 3 times the work it should have done only for computing $\textsc{MatMul}_{r}$, i.e., the protocol is $\alpha$-efficient for $\alpha(t_{\textsc{MatMul}}(n))=3(t_{\textsc{MatMul}}(n)+o(1))$. On the other hand, there is a $\mathsf{Solve}^{*}$ that does essentially only $t_{\textsc{MatMul}}(n)$ work, concluding that the protocol is not an optimal PoUW.
6.5 A Second Instantiation
Our second instantiation of the Encode and Decode operations also relies on the classical random self-reducibility of matrix multiplication, however, here the unpredictability assumption is more novel. In this instantiation we use $r$ as a parameter and explain how it can be instantiated below.
{mdframed}
[ backgroundcolor=codeBackground, linecolor=black, linewidth=0.5pt, roundcorner=5pt, skipabove=10pt, skipbelow=10pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt ]Second Implementation of Encode and Decode
$\textsc{Encode}_{r}(\sigma,A,B)$: $\quad\sslash\quad\sigma\in\{0,1\}^{\lambda},\quad A,B\in{\mathbb{F}}_{q}^{n\times n}$
  1. 1.
    Parse $\sigma=E_{L},E_{R},F_{L},F_{R}$ as four matrices, where $E_{L},F_{L}\in{\mathbb{F}}_{q}^{n\times r}$ and $E_{R},F_{R}\in{\mathbb{F}}_{q}^{r\times n}$.
  2. 2.
    Compute $E=E_{L}\cdot E_{R}\text{ and }F=F_{L}\cdot F_{R}$ (invoking MatMul twice).
  3. 3.
    Output $A^{\prime}=A+E$ and $B^{\prime}=B+F$.
$\textsc{Decode}_{r}(\sigma,C^{\prime})$: $\quad\sslash\quad\sigma\in\{0,1\}^{\lambda},\quad C^{\prime}\in{\mathbb{F}}_{q}^{n\times n}$
  1. 1.
    Parse $\sigma=E_{L},E_{R},F_{L},F_{R}$ as four matrices, where $E_{L},F_{L}\in{\mathbb{F}}_{q}^{n\times r}$ and $E_{R},F_{R}\in{\mathbb{F}}_{q}^{r\times n}$.
  2. 2.
    Compute $C^{\prime\prime}=(A\cdot F_{L})\cdot F_{R}+E_{L}\cdot(E_{R}\cdot(B+F_{L}\cdot F_{R}))$ (invoking MatMul 5 times).
  3. 3.
    Output $C=C^{\prime}-C^{\prime\prime}$.
Algorithm 6.4
It is obvious that completeness holds as
$C^{\prime\prime}=A\cdot F+E\cdot(B+F)$
and therefore
$C^{\prime}-C^{\prime\prime}=(A+E)\cdot(B+F)-A\cdot F+E\cdot(B+F)=A\cdot B.$
(notice that the $E$ and $F$ matrices are consistent between the two procedures). We now discuss transcript unpredictability. Here, $r$ is a parameter and so the transcript of the computation consists of $(n/r)^{3}$ intermediate $r\times r$ matrices. Further, $E$ and $F$ are completely random conditioned on being rank $r$ (see Section 6.5.1 for a proof). The question is whether it is possible to compute the transcript of $\textsc{MatMul}_{r}(A,B)$ without essentially performing the computation as we described in Section 6.1. We formalize this next.
Assumption 6.4.
There is no algorithm to compute all the intermediate values when multiplying two random $n\times n$ rank-$r$ matrices running in time $o(n^{\omega_{r}+1}/r)$.
Recall that the algorithm that we described in Section 6.1 can be used to multiply any two matrices in $O(n^{\omega_{r}+1}/r)$ time, and in particular, to multiply random rank $r$ ones. One can wonder if there is a way to get a speedup by utilizing the low rank property. Suppose that the matrices one wants to multiply are denoted $A$ and $B$, and that the low rank decomposition of them is $A=A_{L}\cdot A_{R}$ and $B=B_{L}\cdot B_{R}$, where $A_{L},B_{L}$ are $n\times r$ and $A_{L},B_{L}$ are $r\times n$. Now, one can generate all the intermediates by computing a partial sum $T_{z}=\sum_{i=1}^{z}A_{R}\cdot B_{L}$ in blocks of size $r$ (so $z$ ranges in $1,\ldots,n/r$), and then multiplying $A_{L}\cdot T_{z}\cdot B_{R}$. Each of the latter can be done by fast rectangular matrix multiplication, costing $n^{\omega_{r}}$, but there are $n/r$ different values of $z$. So, overall, we obtain the same running time of $O(n^{\omega_{r}+1}/r)$, as the general purpose algorithm.
Efficiency of PoUW.
Here, as opposed to all of our previous suggestions, the PoUW is optimal. Indeed, the honest solver $\mathsf{Solve}$ is doing exactly one product of two $n\times n$ matrices, as all of the products in Encode and Decode are products of $n\times n$ and $n\times r$ matrices, that cost $O(n^{2}\cdot r)$. Overall, by a direct analysis,
$\alpha(t_{\textsc{MatMul}}(n))=O(n^{2}+n^{2}\cdot r)+t_{\textsc{MatMul}}(n).$
Note that the last summand is the dominant term. By our assumption on the fastest way to compute the intermediate matrices is by using fast (rectangular) matrix multiplication as described in Section 6.1, there is no algorithm that runs in time $o(t_{\textsc{MatMul}}(n))$, making our protocol an optimal PoUW.
6.5.1 Properties of our Noise Matrices
The method of generating noise matrices described in Algorithm 6.5 results with a uniformly random rank $r$ matrix. Let $\mathcal{E}_{r,n}$ be the set of all $n\times n$ rank $r$ matrices.
Lemma 6.5.
In the random oracle model and assuming that $\sigma$ is unpredictable, then the induced distribution of $E$ and $F$ is uniformly random from $\mathcal{E}_{r,n}$ (with very small statistical error). In particular, with very high probability, every $r\times r$ submatrix of $E$ and $F$ is marginally uniform.
Proof.
We show that $E$ is a uniformly random rank $r$ matrix of dimension $n\times n$. The analog statement for $F$ is proven in the same way.
Let $\mathcal{A}_{n,r}$ denote the set of all $n\times r$ matrices of rank $r$, and let $\mathcal{B}_{r,n}$ denote the set of all $r\times n$ matrices of rank $r$. First, in Claim 6.6, we argue that except with small probability of error, $E_{L}$ is distributed uniformly at random in $\mathcal{A}_{n,r}$. The analog claim holds for $E_{R}$ as well. We thus condition on $E_{L}$ and $E_{R}$ being uniformly random in their respective domains conditioned on being full rank.
Let $\mathcal{M}_{n,n,r}\subset{\mathbb{F}}_{q}^{n\times n}$ denote the set of all $n\times n$ matrices over ${\mathbb{F}}_{q}$ with rank exactly $r$. By construction, the row space of $E$ is the row space of $E_{L}$, and since $E_{L}$ is chosen uniformly from $\mathcal{A}_{n,r}$, the row space of $E$ is uniformly distributed over all $r$-dimensional subspaces of ${\mathbb{F}}_{q}^{n}$. Similarly, the column space of $E$ is the column space of $E_{R}$, and since $E_{R}$ is chosen uniformly from $\mathcal{B}_{r,n}$, the column space of $E$ is uniformly distributed over all $r$-dimensional subspaces of ${\mathbb{F}}_{q}^{n}$. Overall, the row and column spaces of $E$ are independent and uniformly distributed over all possible $r$-dimensional subspaces of ${\mathbb{F}}_{q}^{n}$ and ${\mathbb{F}}_{q}^{n}$, respectively. This means (a) every matrix in $\mathcal{M}_{n,n,r}$ can be produced and moreover each one is equally likely to be produced.
Claim 6.6.
Let $A$ be an $n\times n$ matrix (with $n<n$) whose entries are chosen independently and uniformly from the integer range $[-q,q]$. Then, $A$ is full rank with high probability, namely,
$P(\text{rank}(A)=n)=1-O\left(\frac{1}{q^{n-d+1}}\right).$
Proof.
The matrix $A$ will have rank $n$ if and only if its $n$ columns are linearly independent. For $A$ to be rank-deficient (rank $<d$), there must exist a non-trivial linear combination of the columns that results in the zero vector. This is equivalent to saying that one of the columns lies in the span of the preceding columns.
We can compute the probability that each new column is linearly independent of the previous ones as we proceed through the matrix construction column by column.
The first column of $A$ is non-zero with probability $1-\frac{1}{q^{n}}$, since it must not lie in the zero subspace. Given that the first column is non-zero, the second column is linearly independent of the first with probability $1-\frac{1}{q^{n-1}}$, as it must not lie in the one-dimensional subspace spanned by the first column. More generally, for the $k$-th column to be linearly independent of the previous $k-1$ columns, it must avoid the $(k-1)$-dimensional subspace spanned by those columns. This occurs with probability $1-\frac{1}{q^{n-k+1}}$.
The probability that all $n$ columns are linearly independent is the product of these probabilities, giving that
$P(\text{rank}(A)=n)=\prod_{k=0}^{n-1}\left(1-\frac{1}{q^{n-k}}\right)\geq\prod_{k=0}^{n-1}\exp\left({-\frac{1}{q^{n-k}}}\right)=\exp\left({-\sum_{k=0}^{d-1}\frac{1}{q^{n-k}}}\right),$
where the inequality holds since $\exp^{-x}<1-x$ for $x\in[0,1]$. Since the sum $\sum_{k=0}^{d-1}\frac{1}{q^{n-k}}$ is dominated by its last term, $\frac{1}{q^{n-d+1}}$, we get that
$\sum_{k=0}^{d-1}\frac{1}{q^{n-k}}=\Theta\left(\frac{1}{q^{n-d+1}}\right).$
Therefore, since $1-x+\frac{x^{2}}{2}<e^{-x}$ for $x\in[0,1]$, we conclude that
$P(\text{rank}(A)=d)\geq\exp\left({-\Theta\left(\frac{1}{q^{n-d+1}}\right)}\right)\geq 1-O\left(\frac{1}{q^{n-d+1}}\right).$
7 Using Native AI Workloads for Blockchain Consensus
Artificial Intelligence (AI) tasks, both training and inference phases, can be conceptualized as sequences of matrix multiplication operations interspersed with various other operations such as activation functions, normalizations, and rounding. The matrix multiplications are typically the most computationally intensive parts of any existing networks, with complexities often scaling as cubic in contrast to the square complexity of many of the other operations. The operations performed on these matrices during training and inference can be broken down into a series of matrix multiplications.
Training phase.
During the training phase, the model adjusts its weights based on the input data and the error of its predictions. The key steps involve:
  • Forward Propagation:
    • Linear Transformation: Each layer performs a linear transformation, which is essentially a matrix multiplication between the input data matrix $X$ and the weight matrix $W$. This layer computes $Z=XW$.
    • Activation Function: After the linear transformation, an activation function $f$ (e.g., ReLU, Sigmoid) is applied element-wise to introduce non-linearity.
  • Backward Propagation:
    • Gradient Calculation: The gradients of the loss function with respect to each weight matrix are computed. This involves several matrix multiplications, particularly when applying the chain rule of differentiation through the layers of the network.
    • Weight Update: The weight matrices are updated using the gradients and a learning rate. The update step is also a matrix operation, albeit simpler than multiplication.
Inference phase.
During inference, the model uses the learned weights to make predictions. This primarily involves:
  • Linear Transformation and Activation: Similar to the forward propagation in training, the input data passes through the network layers, each performing a matrix multiplication followed by an activation function.
  • Output Layer: The final layer often includes a matrix multiplication to produce the output predictions.
Reducing cuPOW overhead in AI workloads.
Utilizing the structure and predictability of MatMuls in AI workloads can provide significant dpeedups in practice (i.e., reduction in overhead). We briefly discuss two ideas. First, in several applications, notably inference, the weight matrices are known and are public at the start of the process, and so their hash can be preprocessed, avoiding the need to hash them at every layer “on the fly.” Second, in training applications, the matrices in a given level are typically the output of a product from the previous level (after activations). Thus, in principle, one need not commit to the matrices from scratch in every level, but rather one can utilize a commitment to the very first level and then “recompute” the needed matrices.
Computational complexity and load.
The computational burden in AI models primarily arises from matrix multiplications. For matrices of size $n\times n$ and $n\times m$, the multiplication has a complexity of $nnm$. When $n=n=m$, this simplifies to $O(n^{3})$. Other operations, including activation functions, normalizations, rounding, etc. typically scale with $O(n^{2})$ as they operate element-wise.
Based on experiments and benchmarks we have performed, we estimate that matrix multiplications dominate the computational workload in neural network training. In particular, we find that in typical scenarios at least 50% and often even 70-80% of the total computational load is devoted to matrix multiplication.
Appendix A A Self-Canceling Noise Scheme for high-Rank matrices
The reason we used additive noise (rather than multiplicative) in our PoUW Scheme (Algorithm 6.5) is in order to deal with the case where either $A$ or $B$ are the all-0 matrices (say), in which case any multiplicative operation is futile. Nevertheless, typical useful matrices are never of this form, so one observation is that we may w.l.o.g impose any statistical or algebraic property $\mathcal{P}(A)$ on $A$ (or $B$), so long as it is: (i) satisfied by real-world distributions with high probability $99.9\%$ (or else all bets are off and the protocol rejects); and (ii) efficiently provable and verifiable in sublinear time (ideally, one can implement a SNARK for it in practice). Assuming these conditions are met, restricting $A,B$ to the desired sub-family of matrices (i.e., passing the randomized property-test) opens up the possibility of using alternative—possibly more efficient—noise operators.
Self-Canceling Noise via Pseudorandom Rotations
In particular, we observe that using pseudorandom rotations on high-rank matrices, enables avoiding the decoding step in Algorithm 6.5 of “peeling" the noise altogether. The idea is to randomly rotate $A,B$ using a fast (pseudo-random) orthonormal matrix $R$, e.g., the FastJL Transform [AC09] which allows $O(n^{2}\log n)$-time rotations using the Fast Fourier Transform (FFT), and run the same PoUW scheme from Algorithm 6.5 on the product of
$A^{\prime}:=AR\;,\;B^{\prime}:=R^{\top}B.$
The main point is that, while the output is preserved $A^{\prime}B^{\prime}=ARR^{\top}B=AB$ (no need to “‘peel-off” the noise), this operation completely “scrambles" the partial sums $P_{I,K,J}=\sum_{i\in I,k\in K,j\in J}(AR)_{ik}(R^{\top}B)_{kj}$ of intermediate $r\times r$ tiles ($|I|=|K|=|J|=r$). In particular, assuming the columns of $A$ and rows of $B$ have high rank (say full-rank $n$ for simplicity), the FastJL transofm $R$ guarantees that every (say) $\sqrt{n}\times\sqrt{n}$ tile of $AR$ is a random subspace embedding [CW13], so we may as well assume $A=B=I_{n}$. In this case, computing the partial sums $(RR^{\top})_{I,J,K}\in\mathbb{R}^{\sqrt{n}\times\sqrt{n}}$ seems hard, as it precludes exploiting the global (this time, FFT) structure of $R$ on submatrices – Indeed, it is not known how to multiply arbitrary $r\times r$ submatrices of the FFT matrix by general $x\in\mathbb{R}^{r}$ in $o(r^{2})$ time, unless massive pre-processing is allowed [KU11]. More formally, we conjecture that, for any full-rank $A,B\in\mathbb{R}^{n\times n}$, the amortized cost of computing all partial sums $(ARR^{\top}B)_{I,K,J}$ even conditioned on any (precomputed) function $f(A,B)$, still requires the trivial $\Theta\left(r^{3}\right)$ time (analogue of Assumption 6.4).
Efficient Proof of Proximity for Matrix Rank?
As mentioned earlier, the Pseudorandom-Rotation scheme requires an efficient SNARK (proof of proximity) for testing (or more modestly, distinguishing) high-from-low rank matrices $A,B$ ($\leq k$ vs. $\geq ck$) – In principle, this problem is equivalent to Low-Degree testing of the characteristic polynomial of $A$, which has a fast and practical SNARK ($O(\log^{2}k)$ [RS97, BSBHR18]), alas evaluating $p_{A}(x):=\det(A-x\cdot I)$ at arbitrary points $x$ explicitly has exponentially many terms, hence approximation via sampling seems necessary to make this idea work.
We leave the theoretical and practical exploration of the (pseudo-)random rotation scheme for future work.
Appendix B Proof of Useful Work as a Poisson Process
Remember that a proof of useful work is associated with a function $f\colon\{0,1\}^{n}\to\{0,1\}^{n}$, reflecting the computation being done during the proof of work. We assume a canonical algorithm Alg for solving $f$, and denote $t(n,\mathsf{x})$ the complexity of Alg in solving $\mathsf{x}\in\{0,1\}^{n}$. We denote $t(n)=\max_{\mathsf{x}}f(n,\mathsf{x})$ the worst-case complexity of Alg over all inputs of size $n$.
In a proof of work for blockchains, we require that the event that $\mathsf{Veify}$ outputs 1 follows a Poisson process. A Poisson process is a stochastic process that models the occurrence of random events over time or space, where these events occur independently and at a constant average rate. It is characterized by having exponentially distributed inter-arrival times between events. Bitcoin’s proof of work mechanism is an example of a Poisson process wherein blocks are generated at a constant rate of 10 minutes.
Definition B.1 (Poisson process).
A stream of events $E_{1},E_{2},\ldots$ is a Poisson process with rate $\rho$ if for every point in time $T$, interval $\overline{T}$, and parameter $k$, it holds that
$\Pr\left[\sum_{i=T+1}^{T+\overline{T}}E_{i}=k\right]=\frac{(\rho\overline{T})^{k}e^{-\rho\overline{T}}}{k!}.$
Fact B.2.
Consider a sequence of $n$ independent and identically distributed Bernoulli random variables $E_{1},\ldots,E_{n}$ with parameter $p$. (That is, each $E_{i}$ is 1 with probability $p$ and 0 with probability 00.) The distribution of $\sum_{i=1}^{n}E_{i}$ follows a binomial distribution $B(n,p)$. It is known that the binomial distribution converges towards the Poisson distribution as $n\rightarrow\infty$ and $np$ converges to a finite limit (i.e., if $n$ is sufficiently large and $p$ is sufficiently small).
Definition B.3 (Proof of useful work for blockchains).
Let $\mathsf{Solve}(\sigma,\mathsf{data},\mathsf{x})$ and $\mathsf{Verify}(\sigma,\mathsf{data},\pi)$ be two procedures, as in Section 5, except that we add the $\mathsf{data}$ parameter. These procedures constitute a $(\rho,\alpha)$ proof of work if the following two conditions hold:
  • (a)
    The honest mining process generates a stream of events $E_{1},E_{2},\ldots$ that forms a Poisson process with rate $\rho$.
  • (b)
    For every $T$ and $\overline{T}$, the $(T,\overline{T},\alpha)$-adversarial mining process generates a stream of events $E_{1},E_{2},\ldots$ that forms a Poisson process with rate $\leq\rho$.
{mdframed}
Honest mining process:
  1. 1.
    Initialize $i=1$.
  2. 2.
    Sample a uniformly random $\sigma_{i}$, and choose $\mathsf{x}_{i}$ and $\mathsf{data}_{i}$.
  3. 3.
    Compute $\pi_{i}=\mathsf{Solve}(\sigma_{i},\mathsf{data}_{i},\mathsf{x}_{i})$.
  4. 4.
    Compute $E_{i}=\mathsf{Verify}(\sigma_{i},\mathsf{data}_{i},\pi_{i})$.
  5. 5.
    Increase $i$ by 1 and return to step 2.
$(T,\overline{T},\alpha)$-adversarial mining process:
  1. 1.
    Run the honest process for $T$ steps resulting with $E_{1},\ldots,E_{T}$ and $\{(\sigma_{i},\mathsf{x}_{i},\mathsf{data}_{i},\pi_{i})\}_{i\in[T]}$.
  2. 2.
    Given all of the above, initialize (in polynomial time) a time $\alpha(n)\cdot(t(n)\cdot\overline{T})$ adversary. Upon a set of uniformly random $\sigma_{T+1},\ldots,\sigma_{T+\overline{T}}$, run it to generate $\pi_{T+1},\ldots,\pi_{T+\overline{T}}$. Compute $E_{i}=\mathsf{Verify}(\sigma_{i},\mathsf{data}_{i},\tilde{\pi}_{i})$ for $T<i\leq T+\overline{T}$.
  3. 3.
    Continue running the honest process with $i=T+\overline{T}+1$.
Bitcoin’s proof of work.
To see that Bitcoin’s PoW process forms a Poisson process with rate $\rho$, fix a point in time $t$ and a polynomially long interval $T$. A polynomial number of random oracle evaluations are independent of one another (and of any preprocessing). Thus, the number of successful random oracle evaluations (block discoveries) follows a Binomial distribution: $X\sim\text{Binomial}(n,p)$ with $n=R\cdot T$ and $p=\kappa/2^{\lambda}$, where $R$ is the total hash rate (hashes per second) of the network and $\kappa$ is a difficulty parameter. By Fact B.2, for large $n$ and small $p$, the Binomial distribution approximates a Poisson distribution with parameter $\rho=n\cdot p=(R\cdot T\cdot\kappa)/2^{\lambda}$. The value of $\kappa$ is adjusted so that $\rho=1/600$, implying a block every 10 minutes, in expectation. Since we are using a random oracle to model the hash function, it is also the case that iterated hashes is the best strategy for a malicious solver in the PoW, implying that there is no way to obtain any speedup.
Bitcoin from our PoUW.
We argue that a PoUW for blockchains, as in Definition B.3, along with standard properties of the cryptographic hash function, can be used to establish the security of the Bitcoin backbone protocol [GKL24]. More precisely, PoUW can be used to instantiate the same functionality as that of Bitcoin, i.e., a public ledger, as defined in [GKL24]. The proof follows the one presented in [GKP20], where we instantiate signatures of work with PoUW. We refer to [GKP20] for the proof.